OpenStudy (anonymous):

|x-8/2|=2 could some explain if i have to do two equation one thats the original and the other everything thats neg or positive switch it

OpenStudy (anonymous):

yes you have to solve two equations

OpenStudy (anonymous):

you don't know if \(\frac{x-8}{2}\) is positive or negative it if is positive, then \(\frac{x-8}{2}=2\) and if it negative, then \(\frac{x-8}{2}=-2\)

OpenStudy (anonymous):

so you dont change any signs in the |box| only outside